OpenStudy (anonymous):

An angle measure of 91 degrees is equivalent to ____ radians. Round your answer to the nearest hundredth when necessary.

OpenStudy (solomonzelman):

180 degrees = π so 91/180= 0.51 π

OpenStudy (anonymous):

so it's 0.51? would it just be 0.5 if i round it to the nearest hundredth?

OpenStudy (solomonzelman):

your question asked you to round to nearest hundredth.

OpenStudy (solomonzelman):

0.5 ---> rounded to nearest 10th 0.51 ---> rounded to nearest 100th

OpenStudy (solomonzelman):

@lowcard2, do you need more help on this problem?

OpenStudy (anonymous):

no im good thank you! i understand now cx

OpenStudy (solomonzelman):

Anytime!!!

Directrix (directrix):

@lowcard2 @SolomonZelman >> so it's 0.51 ?? 0.51 π radians approximates to 1.60 radians

OpenStudy (solomonzelman):

@Directrix, I don;t think so.

OpenStudy (solomonzelman):

I plugged into my calc, 91 / 180, and got 0.5055555555555...............

OpenStudy (anonymous):

when i put it in directrix do i put it in as 1.6 or 1.60 ?

Directrix (directrix):

You have to multiply by pi. One radian is roughly 57 degrees so 91 degrees has to be larger than 1.

OpenStudy (solomonzelman):

I see, I am sorry!

OpenStudy (anonymous):

which one is more accurate? 1.6 or 1.60?

Directrix (directrix):

To the nearest hundreth, 1.60 because it cranks out as 1.60221225333

OpenStudy (anonymous):

oh ok thx !cx

Directrix (directrix):

@lowcard2 Hold on while I crank out the problem from the beginning.

Directrix (directrix):

91/180 = x/pi where x is the measure of the angle 180*x = 91*pi x = (91*pi)/180 x = 1.59

Directrix (directrix):

@lowcard2 Do your instructions tell you to use 3.14 for pi?

Directrix (directrix):

@lowcard2

OpenStudy (anonymous):

im not sure thats all it says is up there ^^

Directrix (directrix):

3.14 as an approx for pi gives this: x = 1.59 @lowcard2